Home‎ > ‎Reports‎ > ‎Crew110a‎ > ‎

day06


Crew #110A Log Book for January 5, 2012

Photos of the day





Commander's Check-In Report

Date and time: 5 January 2012, 19:33
Written by: Julie Mason, Crew 110, Commander

Crew Physical Status: Great

Time Departed/Returned from EVA: EVA 1 12:00-13:30, EVA 2 14:00-16:30;

Telescope EVA (Last Night): 19:30 – 05:00

Brief Narrative of Field Mission Results:

Last night four crew members imaged Polaris and Sirius at the Musk Mars Desert Observatory at one hour time increments over the course of the night. We also imaged Venus between increments and observed Saturn around 5:00am. The crew processed images using RegiStax 6 today.

Crew members continued to exercise in the habitat while not on EVA, as part of a habitat exercise study.

During EVA 1, three crew members worked on a mobility study, hiking and throwing a Frisbee in between taking outreach photos.

During EVA 2, three crew members finished taking the last few photos for the 3D image of the habitat and then participated in hiking as part of a mobility study.

Engineering/Hab Maintenance: Spark on stove

Report Transmission Schedule:

Plans for Tomorrow: Two more 3D models will be made tomorrow. EVA 1— Three crew members will take images for a 3D model of a hill in the area. EVA 2— The other three crew members will take images for another 3D model. EVA 3—Observatory, Atmospheric Study (if the sky is clear)

Inventory: Need more oil for the ATV’s

Support Requested: None

Miscellaneous: The crew watched a Saturn and Mars episode of The Universe, a History Channel series, last night.

Commander's Report

Date and time: 5 January 2012, 19:33
Written by Julie Mason, Crew 110A, Commander

Crew members exercised early this morning and got a later start than usual for both EVA’s. Aaron only needed a few more photos for his main 3D image. The other two EVA’s focused on space suit mobility and outreach.

Data is looking good for the crew’s atmospheric study. More data will be collected tonight. Tomorrow, the crew will focus on collecting more images for two more 3D models.


3D Model Project Report

Date and time: 5 January 2012, 17:40
Written by Aaron Olson, Mission Specialist

During the sixth day of Crew 110A’s stay at the MDRS site, Mark, and I took images of areas that didn't turn out well in our original 3D model. I will upload these photos and prepare an updated model with them. We will be traveling to two locations tomorrow to image areas much further away from the habitat. I am hopeful that I will have a completed and complete model of the MDRS habitat area before our crew leaves this coming Saturday.

EVA 7 Report

Date: 5 January 2012
Time: 18:35
Written By: Lyndsey Bankers, Crew 110a, Mission Specialist

Time of EVA: 12:00-13:30

Crew Members: Julie Mason, Lyndsey Bankers, and Will Yu

Site Location: Near the Habitat, about 2 miles North of the Habitat in
an open area

Transit Mode: ATV’s

Objectives: To utilize hiking and a frisbee to study mobility in the
space suits.

Results: Frisbee showed that the space suits limit the mobility of the
arms when raising them above your head. They also hinder running and
make it hard to see different angles that you may want to see when
throwing or catching. Hiking showed that the suits do not hinder the
mobility of the legs, but heavy breathing after hiking up an incline
will fog up the helmets quickly.

Lessons Learned: Frisbee was a bit challenging at first, but it was
easy to adapt and playing became less challenging after a while.


EVA 8 Report

Date: 5 January 2012
Time: 18:35
Written By: Mark Ruff, Crew 110a, Mission Specialist/Journalist.

Time of EVA: 14:00-16:30

Crew Members: Aaron Olsen, Sam Marron, Mark Ruff

Site Location: Hab Area, Hills near the end of Lowell Pass

Transit Mode: ATV’s

Objectives: To capture additional images for 3D imaging study and experiment with hiking in the suits.

Results: The Images captured will serve to add detail to the current 3D model of the area surrounding the hab. Hiking in hills near the end of the Lowell Pass proved viable despite limitations imposed by the suits themselves.

Lessons Learned: Helmet fog is often a limiting factor in the utility of a participant on EVA.

EVA 9 Report

Date: 5 January 2012
Written by: Sam Marron, Crew 110a

Time of EVA: 19:30-5:00

Crew Members: Sam Marron, Will Yu, Lyndsey Bankers, Julie Maso
n

Site Location: Observatory

Transit Mode: Walking through observatory tunnel

Objectives: To video record two stars at hour intervals throughout the night

Results: 3 minute videos of Polaris and Sirius on the hour from 8:00pm until 5:00am. Post processing and image analysis will be used to estimate the amount of extinguishing material in the atmosphere

Lessons Learned: Each star takes roughly 5 minutes to find on the telescope and 5 minutes to set up the video. Occasionally stars can be hard to find, taking upwards of twenty minutes. The initial data shows a good correlation between angle off of the horizon and light intensity, which will provide a good estimation of the extinguishing material in the atmosphere. Difficult to work long hours through the night after participating in an EVA during the day, and vice versa.

Engineering Report

Date and time: 05 January, 2012, 18:30 MST

Written by Will Yu, Crew 110A, Chief Engineer.

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): .75
Propane Tank (Full/0.75/0.5/0.25/Empty): .50
Gasoline Tank (Full/0.75/0.5/0.25/Empty): .50

Oil:
ATV Oil Quantity (# of Quarts in storage): 0
Car Oil Quantity (# of Quarts in storage): N/A

Water Status:
Outside Potable Water Tank Level (inches from bottom): 14
Trailer Potable Water Tank Level (inches from bottom): 26
Hab Potable Water Tank Level (inches from bottom): 12
Potable Water Meter Reading: 48633.7
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): YES
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): NO
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: 125488.9
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: YES
Oil Checked: YES
Fuel Consumed: .5 gal
Tire Status: Good - 4psi
Today's Run Time: 00:40
Season's Run Time: 46:10
Comments:

ATV 2 (Spirit):
Used: YES
Oil Checked: YES
Fuel Consumed: .5 gal
Tire Status: Good - 4psi
Today's Run Time: 00:40
Season's Run Time: 37:33
Comments: Oil level at half full

ATV 3 (Viking I):
Used: YES
Oil Checked: YES
Fuel Consumed: .5 gal
Tire Status: Good
Today's Run Time: 00:40
Season's Run Time: 42:55
Comments: Has broken tail light cover, tail light still works. Back brakes didn’t engage due to loose brake cable. Re-tighten and works fine.

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): Upstairs furnace not turned on since new downstairs furnace was enabled
Actual temperature upstairs (Farenheit): 70 F
Thermostat setting downstairs (Farenheit): Downstairs furnace was on during 01:00 – 06:00
Actual temperature downstairs (Farenheit): 67 F

Computers/Networking Infrastructure:
Notes/comments on internet/computers?:

General Engineering Notes/Comments:
~ Helmet 5 had an existing crack which has no propagated a little further
~ Spark on the stove no longer lights the burners. Manual lighting of the stove required
~ Need ATV oil for available for storage
~ ATV3 has broken tail light cover, tail light still works. Back brakes didn’t engage due to loose brake cable. Re-tighten and works fine.

Journalist Report

Date and time: 5 January 2012 19:00

Written by: Mark Ruff, Crew 110a, Mission Specialist/ Journalist

Only two days left on Mars. Today, Lyndsey, Julie, and Will took the first Eva and had a little fun really putting the suits through their paces. They played some Frisbee and went for a hike, to see just how limiting the suits are to that kind of physical activity. Meanwhile, Aaron, Sam, and I worked on various projects, and made lunch.

After the first EVA team got back, we went out to get the last few pictures for the 3D imaging study, after which, we took the ATV’s out to an area with some really cool rock formations, and got some great photos. It stands to reason that for a Mars program to continue beyond the first few “flags and footprints” missions, they will need to capture the attention of the public back home. One of the best ways to do this would be through the images that are brought back from the red planet. Having found some great scenery, we snapped photos while Sam made bad rock formation puns. Not a bad afternoon.

Tomorrow is our last day on Mars. It’s something we all seem to be regarding with anticipation and at least a twinge of anxiety. It will be nice to get back to the luxuries of life back home, but at the same time, this is the kind of experience you’re never quite ready to leave behind.


Chef's Report


Date: 5 January 2011

Written by Aaron Olson, Crew 110A


1a) Today was a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? If so, what was special about today and what food(s) or activities were prepared to mark the celebration?
No


2) 2 boxes of potatoes au gratin, freeze dried ground beef and diced beef, peanut butter and chocolate bread, 2-layer brownies

3) half a loaf of peanut butter and chocolate bread

4) potatoes were baked in the oven, beef was re-hydrated over the stove range, bread was made in bread maker, brownies were baked
5) The peanut butter and chocolate bread didn't taste much like chocolate at all and the beef was fairly good.